3. Compute the GCD of the following pairs of numbers.
a. 16 and 514 b. 21 and 134 c. 91 and 17
Answer #1

using prime factorizationa
1)16 and 514
16=2*2*2*2
514 = 2*257
Common factor =2 which is the GCD

2)21 and 134
21 = 7*3
134 = 2*67
No common factor, GCD = 1

3)91 and 17
17 = 17
91 = 13*7
No common factor, GCD = 1